I think I read a burette wrong --- in reading a burette should the final reading be higher or lower than the initial reading. I got a final reading of 27.49mL and an initial reading of 47.75mL. If I subtract the higher from the lower a will get a negative number?
1
Expert's answer
2015-10-28T04:11:46-0400
In this case you are interested in the difference volume, so it doesn`t actually matter what you put first and second in the subtraction, you can use resulting number anyway by just removing (-). However, more correct is of course to subtract smaller number from the bigger, like this:
V` = V(start) – V(end) = 47.75 - 27.49 = 20.26 (ml)
Also don`t forget to check meniscus correctly depending on the liquid you measure.